OpenStudy (anonymous):

Let's say I have the equation c = A*P^2 + Q Do I have to solve it for q by minusing q from both sides first

OpenStudy (anonymous):

Or is there another way to solve q

OpenStudy (unklerhaukus):

leave the Q where it is, minus A*P^2 from both sides

OpenStudy (unklerhaukus):

\[c = A*P^2 + Q\] \[c-A*P^2 = Q\]

Parth (parthkohli):

c - q = a*p^2 So, c = (c - q) + q c = c - q + q c = c We can't solve for c because there is no constant.

OpenStudy (anonymous):

I thought u have to divide A*P^2????